OpenStudy (anonymous):

2xraiseto two minus 3x minus nine 6xraise to two minus 27x minus 15 4xraise to two 2x minus 6 5xraise to two minus 18x plus 14 .. please answer .. i need it on my assignment .. or maybe teach me how to get the answer ..

OpenStudy (anonymous):

1- 2x^(2-3x-9) = 2- 6x^(2-27x-15) = 3- 4x^(2-2x-6) = 4- 5x^(2-18x+14) = What are you trying to find out here?